40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 1 12:12:35 2013, gautam, Update, CDS, New model for endtable PZTs C1ASX_GDS_TP.png
    Reply  Thu Aug 1 13:09:17 2013, gautam, Update, CDS, New model for endtable PZTs-procedure 
    Reply  Thu Aug 1 15:06:59 2013, jamie, Update, CDS, New model for endtable PZTs 
       Reply  Thu Aug 1 15:28:44 2013, gautam, Update, CDS, New model for endtable PZTs-problem solved 
    Reply  Thu Aug 1 20:58:56 2013, gautam, Update, CDS, New model for endtable PZTs-MEDM Screens setup MAIN_SCREEN.pdfMATRICES.pdfLOCKINS.pdfCONTROL_FILTERS.pdf
       Reply  Mon Aug 5 18:18:32 2013, gautam, Update, CDS, Choosing LO Amplitudes and Frequencies  GTR_Power_Spectrum.pdf
          Reply  Mon Aug 5 19:10:01 2013, Koji, Update, CDS, Choosing LO Amplitudes and Frequencies  
             Reply  Tue Aug 6 16:36:51 2013, gautam, Update, CDS, Choosing LO Amplitudes and Frequencies-revised powerSpec0806.pdf
                Reply  Wed Aug 7 23:40:49 2013, gautam, Update, CDS, X-End Green ASS - A first update time-series.pdf
                   Reply  Sat Aug 10 05:53:51 2013, gautam, Update, CDS, X-End Green ASS - Roundup New_ASX_MEDM_MAIN.pdf
                      Reply  Mon Aug 12 12:57:59 2013, Jenne, Update, CDS, X-End Green ASS - Roundup 
                         Reply  Mon Aug 12 13:30:33 2013, Jamie, Update, CDS, X-End Green ASS - Roundup 
Message ID: 8972     Entry time: Tue Aug 6 16:36:51 2013     In reply to: 8968     Reply to this: 8983
Author: gautam 
Type: Update 
Category: CDS 
Subject: Choosing LO Amplitudes and Frequencies-revised 

I redid the power spectrum measurement for the X-arm green transmission after aligning the arm to green using the ITMX/ETMX Pitch and Yaw sliders on IFOalign.

The Y-axis now reflects the relative intensity noise (RIN), which I obtained by taking the average value of the X-arm green transmission using tdsavg. Based on this measurement, I have now picked four new frequencies at which to try and modulate the PZT mirrors: 10, 19, 34 and 39 Hz. Bandpass filters in the LIA stage have been appropriately modified. 

Power Spectrum:

powerSpec0806.pdf

ELOG V3.1.3-